se the table below for Problems 22 - 25. The debt and equity columns describe different possible capital structures A-G for Company XYZ. The Method 1-3 columns indicate the WACC for capital structures A-G and are consistent with the Modigliani and Miller Propositions Structure Total Assets 8000 4000 1200 1000 800 600 400 Equity 400 400 400 400 400 400 400 Method 1 16.13% Method 2 5.49% 6.28% 13.20% 12.24% 11.00% 11.20% 13.00%) Method 3 13.00% 13.00% 13.00% 13.00% 13.00% 13.00% 13.00% 972% 9.60% 10.73% 13.00%) 22. The market risk premium is 10% and the risk free rate is 4.0%. Using the CAPM, calculat e the (unlevered) beta for Company XYZ: A, 077 B. 0.90 C. 1.00 D. 1.27 E. 1.30 23. Which method(s) imply that the tax shield from debt can at least sometimes reduce wACC? A. Method 1 B. Method 2 C. Method 3 D. Method 1& Method 2 E. Method 1 & Method3 24. Assume Company XYZs currently has $1000 in total assets and the higher the firms debt level, the higher the risk of bankruptcy. If the CFO is looking to maximize firm value, Company XYX should: A. borrow an additional $400 in debt B. repay or retire $400 in debt C. borrow an additional $200 in debt D. repay or retire $200 in debt E. make no changes to the capital structure 25. Assume Company XYZs currently has capital structure C under Method 1 where the WACC is 13.20% based on an effective tax rate of 30% and a cost of debt of 18%. There are 200 shares currently outstanding. The company just reported (TO) FCFF of 300 and FCFE of 250. All cash flows are expected to grow at 20% for each of the next three years, then at 6% beginning in year four (T-4) in perpetuity (or indefinitely). If you were an activist investor trying to buy gain control of Company XYZ, what is the most you would pay per share for the hostile takeover? A. $6.97 B. $9.62 C. $11.55 D. $15.49 E. $31.36
Answer #1

Question 22:

When we talk about unlevered beta, this means the capital structure should have no debt at all. This is consistent with capital structure G where out of 400 assets, entire 400 is equity. So, under this case the cost of equity is 13%

So, as per CAPM:

Re = Rf + beta * Market risk premium

13% = 4% + beta * 10%

13 =4 + 10*beta

beta = 9/10 =0.9

Unlevered beta = 0.9 (Option B)

Note: We have answered one full question. Please note that only one multiple choice question can be answered at a time. Kindly post each separately for experts to answer